Suppose that weights of 5th grade elementary girls are normally distributed with mean µ = 64...

  1. Suppose that weights of 5th grade elementary girls are normally distributed with mean µ = 64 lbs and standard deviation σ = 6.6 lbs. Find the weight that corresponds to P42 and interpret this measure of position.
Answer #1

solution

Using standard normal table,

P(Z < z) = 42%

= P(Z < z) = 0.42  

= P(Z < -0.20) = 0.42

z = -0.20 Using standard normal z table,

Using z-score formula  

x= z * +

x= -0.20* 6.6+64

x= 62.68

x=63